OpenStudy (anonymous):

Brian is 2 years older then Bob who is 7 years older then Mark, Their aged combined is 61, find the age of each person?

OpenStudy (anonymous):

so let x become mark's age, x+7 become bob's, x+9 become brain's

OpenStudy (anonymous):

x+x+7+x+9=61

OpenStudy (anonymous):

3x+16=61

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

so mark=15, bob 15+7=22, brain 15+7+2=24

OpenStudy (anonymous):

great

OpenStudy (anonymous):

thx...

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

thank you